A 5.0-kg block slides from an initial speed of 8.0
Answer #1

The block stopped due to retardation imposed by friction,

Work done to overcome friction = Heat generated

Initial energy = 5 (9.8) 12 sin 15 + 1/2 ( 5)(8)^2 = 172.68 J

Final Energy = 0

Heat generated = 172.68 J

\DeltaS ( entropy chnage0 = 172.68/ (300 K) = 0.5756 J/K